# Nearly geodesic surfaces are filling

### v2

@@ Abstract
Let $M$ be a closed hyperbolic $3$-manifold. A homotopy class $[S]$ of surfaces in $M$ is filling if for any $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components that are contractible in $M$. We prove that there exist $\epsilon_{0},g_{0}>0$ such that every homotopy class of $(1+\epsilon)$-quasi-Fuchsian surfaces with $0<\epsilon\leq\epsilon_{0}$ or totally geodesic surfaces of genus $\geq g_{0}$ in $M$ is filling. As a corollary, except for at most finitely many totally geodesic surfaces, embedded incompressible quasi-Fuchsian surfaces in $M$ have constants bounded below by $1+\epsilon_{0}$. This also gives a gap theorem for embedded minimal surfaces. Each of these surfaces separates any pair of distinct points at $\partial\widetilde{M}$. Crucial tools include the rigidity results of Mozes-Shah, Ratner, and Shah. This work is inspired by a question of Wu and Xue whether random geodesics on random hyperbolic surfaces are filling.
@@ 1 Introduction
Let $M$ be a closed hyperbolic $3$-manifold and $\widetilde{M}$ the universal cover. A surface $S$ in $M$ is $(1+\epsilon)$-quasi-Fuchsian if its limit set $\partial\widetilde{S}\subset\partial\widetilde{M}$ is a $(1+\epsilon)$-quasicircle (but not a $(1+\epsilon^{\prime})$-quasicircle for any $\epsilon^{\prime}<\epsilon$). A homotopy class $[S]$ is filling if for every $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components contractible in $M$. Rubinstein-Sageev [RS99] define $[S]$ to be strongly filling if for any $p\neq q\in\partial\widetilde{M}$, there is an $\widetilde{S}$ covering $S$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S}$. The genus of a totally geodesic surface is the smallest genus of an orientable surface in its commensurability class. In this paper, we prove
@@
Let $M$ be a closed hyperbolic $3$-manifold. There exist $\epsilon_{0},g_{0}>0$ such that every homotopy class of $(1+\epsilon)$-quasi-Fuchsian surfaces with $0<\epsilon\leq\epsilon_{0}$ or totally geodesic surfaces of genus $\geq g_{0}$ is strongly filling.
Kahn and Marković [KM12] prove that $M$ has infinitely many $(1+\epsilon)$-quasi-Fuchsian surfaces for $\epsilon<\epsilon_{0}$. Moreover, the proof of Theorem 1.1 shows that all but finitely many totally geodesic hypersurfaces of a closed hyperbolic manifold are strongly filling. An immersed surface $S\subset M$ is $\pi_{1}$-injective if the induced homomorphism $\pi_{1}(S)\rightarrow\pi_{1}(M)$ is injective. For the rest of the paper, we set $\epsilon_{0}$ and $g_{0}$ as in Theorem 1.1.
@@
Let $M$ be a closed hyperbolic $3$-manifold. The set of quasi-Fuchsian $\pi_{1}$-injective embedded surfaces in $M$ is decomposed into
@@
    at most finitely many totally geodesic surfaces of genus $g_{0}$, and
@@
The resolution of the tameness conjecture [Ago04, CG06] implies that a $\pi_{1}$-injective surface in $M$ is either quasi-Fuchsian or virtually fibered (lifts to a fiber of a finite cover that is a surface bundle over $S^{1}$). The supremum principal curvature of a fibered minimal surface $F$ is >1>1 by [Uhl83, HL21] and the Hausdorff dimension of the limit set of $F$ is $2$. Thus 1.2 implies that any sequence of $\pi_{1}$-injective surfaces in $M$ is eventually self-intersecting or uniformly away from being totally geodesic.
@@
Let $M$ be a closed hyperbolic $3$-manifold. If $S\hookrightarrow M$ is an embedded minimal surface, then $S$ is either a totally geodesic surface of genus $g_{0}$ or a quasi-Fuchsian surface of supreme principal curvature $\geq\frac{\epsilon_{0}}{\epsilon_{0}+2}$.
@@
We say $[S]$ in $M$ separates two distinct points $p,q\in\partial\widetilde{M}$ if $S^{\prime}\in[S]$ has a cover $\widetilde{S^{\prime}}\subset\widetilde{M}$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S^{\prime}}$. Kahn-Marković [KM12] constructs a collection of nearly geodesic surfaces which is dense in the Grassmannian $2$-plane bundle $\mathcal{G}_{2}M$ and Bergeron-Wise [BW12, Corollary 4.2] points out that these nearly geodesic surfaces satisfy a boundary criterion.
Theorem 1.4 (KM12).
@@
Let $M$ be a closed hyperbolic $3$-manifold. Every pair of distinct points $p,q\in\partial\widetilde{M}$ is separated by some quasi-Fuchsian $[S]$ in $M$.
This theorem, along with other important results of Agol, Bergeron-Wise, Haglund-Wise, Sageev, and Wise, has important consequences in geometric group theory and virtual properties of $3$-manifolds, see Section 3.6, [Ago13, HW08, Sag95, Wis21], and the references therein. The boundary criterion of Bergeron-Wise provides a motivation to study strongly filling surfaces .
@@
If $[S]$ is strongly filling, then for any $^{\prime}\in[S], SS^{\prime}$ intersects every homotopically nontrivial closed curve in $M$, which implies that $[S]$ is filling. By 2.7, a least area representative of a filling homotopy class actually cuts the ambient manifold into topological balls. In Section 3.6 we point out that a strongly filling $[S]$ gives a cubulation with a single orbit of hyperplanes (whose stabilizer is essential). Non commensurable strongly filling surfaces give rise to cubulations that are distinct under the isometric actions by $\pi_{1}(M)$ and thus their existence answers a question of Fioravanti-Hagen on [FH21, p. 879].
@@
Figure 1: The filling nature of nearly geodesic surfaces in a closed hyperbolic $3$-manifold (each dot represents a $(1+\epsilon)$-quasi-Fuchsian surface with $\epsilon\leq\epsilon_{0}$).
Before talking about the proof of Theorem 1.1, we summarize it in Figure 1, which illustrates the idea that a certain neighborhood of totally geodesic surfaces of genus $\geq g_{0}$ in the space of all $\pi_{1}$-injective surfaces consists of filling surfaces. In various arguments, it is helpful for the author to view transverse intersections and filling as an “open” property of $\mathcal{G}_{2}M$ and of the space of $\pi_{1}$-injective surfaces in $M$. This figure is inspired by [KMS23, Figure 1] and justified in Section 2.1.
Two $\pi_{1}$-injective surfaces $S_{1},S_{2}\subset M$ are commensurable if there exists a $g\in\pi_{1}(M)$ such that $g\pi_{1}(S_{1})g^{-1}\cap\pi_{1}(S_{2})$ is of finite index in both $\pi_{1}(S_{1})$ and $\pi_{1}(S_{2})$. A commensurability class of surfaces is filling if each of its homotopy classes is filling. We always assume surfaces are orientable and $\pi_{1}$-injective unless stated otherwise. A surface is Fuchsian or totally geodesic if any geodesic on $S$ with respect to its induced Riemannian metric is also a geodesic on $M$. Commensurable Fuchsian surfaces correspond to the same $PSL_{2}(\mathbb{R})$-orbit. A sequence of $(1+\epsilon)$-quasi-Fuchsian minimal surfaces is asymptotically Fuchsian if $\epsilon\rightarrow 0$. The proof of Theorem 1.1 uses Theorem 1.6 and 3.15.
@@
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{j}$ is a sequence of asymptotically Fuchsian surfaces in distinct commensurability classes, then the Grassmann $2$-plane bundle $\mathcal{G}_{2}S_{i}\rightarrow\mathcal{G}_{2}M$ in the Hausdorff metric.
Remark 1.7.
Independently, Fernando Al Assal and Ben Lowe [AL25] obtain Theorem 1.6 for finite-volume hyperbolic manifolds with different perspectives and applications. Al Assal [AA23] shows that the measure limits of $\mathcal{G}_{2}S_{i}$ achieve every convex combination of $SL_{2}(\mathbb{R})$-invariant ergodic measures.
Vladimir Marković suggested the “one versus uncountably many planes duality” for the limiting behavior of asymptotically Fuchsian surfaces to the author, which led to Theorem 1.6. 3.15 then establishes filling by the following argument. Suppose there exist nontrivial $\gamma_{i}]\in\pi_{1}(M-S_{i})$. Then a subsequence of $(x_{i},v_{i})$ in the unit tangent bundle $T^{1}\gamma_{i}$ of the geodesic $\gamma_{i}$ converges to $(x,v)$ tangent to a geodesic $\gamma$. There exist $(p_{j},\Pi_{j})\in\mathcal{G}_{2}S_{j}\rightarrow(x,\Pi)\perp(x,v)$. Since $\epsilon_{j}\rightarrow 0$, large $S_{j}$ are nearly orthogonal to $\gamma_{j}$, and then $\partial\tilde{S}_{j}$ separates $\partial\widetilde{\gamma_{j}}$ which means $\pi_{1}(M-S_{j})$ cannot contain $\gamma_{j}$. We need to use [Sep16] which controls the norms of the second fundamental forms of $\tilde{S}_{j}$ by $\epsilon_{j}$ and [Smi07] to prove $\tilde{S}_{j}$ converge to a totally geodesic plane $\widetilde{P}$.
A geodesic plane in $M$ is the image of a totally geodesic plane $\widetilde{P}\subset\mathbb{H}^{3}\rightarrow M$. The proof of Theorem 1.6 uses crucially [Rat91, Sha91a] which imply a geodesic plane in $M$ is either dense or a closed Fuchsian surface, and uses [MS95] which implies that a sequence of Fuchsian surfaces in distinct commensurability classes equidistribute (become uniformly dense in $\mathcal{G}_{2}M$). Since $S_{i}$ are minimal, their Grassmann bundles are connected closed sets, and their Hausdorff limit is a connected closed subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. Any finite union of the Grassmann bundles over Fuchsian surfaces is a disconnected closed set. Thus, $\mathcal{G}_{2}S$, which is closed and $PSL_{2}(\mathbb{R})$-invariant, must contain a dense plane in $\mathcal{G}_{2}M$.
Theorem 1.1 is inspired by Theorem 1.8. Given a Fuchsian $S$, let $\mathbf{g}_{c}(S)$ be the smallest genus of an orientable surface in $M$ commensurable with $S$.
@@
Let $M$ be a closed hyperbolic $3$-manifold. Given a sequence of Fuchsian surfaces $S_{i}\hookrightarrow M$ with $\mathbf{g}_{c}(S_{i})\rightarrow\infty$, the maximal diameter of components of $M-S_{i}$ tends to 0 and the number of components tends to infinity.
The proof of Theorem 1.8 actually shows that a sequence of totally geodesic hypersurfaces cuts the ambient closed hyperbolic $n$-manifold ($n\geq 3$) into components whose diameters tend to 0. Since $M$ is aspherical, a component of MSM-S of diameter $2\textnormal{inj}(M)$ is necessarily contractible. By [BFMS21] (for $n\geq 3$) and [MM22] (for $n=3$), if $M$ contains infinitely totally geodesic hypersurfaces, then $M$ is arithmetic.
An outline of the proof of Theorem 1.8 is the following. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Suppose there exist components $C_{i}\subset M-S_{i}$ with $\textnormal{diam}(C_{i})\geq 2R$. Since $\pi^{-1}(C_{i})\subset\mathbb{H}^{3}$ consists of convex components, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length =2R=2R. Up to a subsequence, $\gamma_{i}\rightarrow$ a geodesic segment $\gamma$ of length $2R$. 3.7 proves that the subset of $\mathcal{G}_{2}M$ whose points are tangent to the planes transversally intersecting the central half of $\gamma$ with uniformly bounded angles contains an open set of $\mathcal{G}_{2}M$ of uniformly positive measure (depending only on $R$). Thus eventually $S_{i}$ must intersect the central half of $\gamma$ with uniformly bounded angles. This shows that $S_{i}$ must transversally intersect $\gamma_{i}$ for large $i$, a contradiction. If the maximal diameter of every component of $M-S_{i}$ is less than $2\textnormal{inj}(M)$, then $S_{i}$ is filling. [LM23] points out the rate of equidistribution and hence $g_{0}$ depends only on the rate of mixing, $\textnormal{Vol}(M)$, and the injectivity radius. 3.18 proves that a filling Fuchsian surface necessarily has some transverse intersections with all geodesics in $M$, which we are uncertain to hold for filling least area non Fuchsian surfaces.
One way to visualize Theorem 1.8 and its transition to Theorem 1.1 is the following. Let $D$ be a convex polyhedron fundamental domain for $M$ in $\mathbb{H}^{3}$. The generators of $\pi_{1}\rightarrow\Gamma\leq PSL_{2}(\mathbb{C})$ are given by finitely many side-pairings of $\partial D$, i.e., $\{g\in\Gamma:g\overline{D}\cap\overline{D}\subset\partial\overline{D}\textnormal{ and}\neq\emptyset\}$. The preimages of large filling Fuchsian surface give rise to some geodesic planes that cut $D$ and $\partial D$ into components so that any geodesic that passes through $D$ must intersect one of the planes transversally.
In addition to being an $SL_{2}(\mathbb{R})$-orbit in $\mathbb{H}^{3}$, a geodesic plane can be defined as a submanifold with principal curvatures 0. A slight perturbation to these geodesic planes gives rise to minimal planes with small principal curvatures and limit sets with small quasi-Fuchsian constants. Since the polyhedron fundamental domain is compact, finitely many evenly distributed almost-geodesic planes remain candidates for filling. In other words, although the proof of Theorem 1.8 relies on the rigidity of $SL_{2}(\mathbb{R})$-orbits, the topological filling nature persists for nearby surfaces which are not homogeneous (i.e., orbits of Lie subgroups of $SL_{2}(\mathbb{C})$).
A stronger notion of separation than filling is ubiquitous. A collection of immersed surfaces in a hyperbolic 33–manifold $M$ is ubiquitous if, for any pair of disjoint and nonasymptotic planes $\Pi,\Pi^{\prime}\subset\mathbb{H}^{3}$, some surface in the collection has an embedded preimage $\tilde{S}\subset\mathbb{H}^{3}$ that separates $\Pi$ from $\Pi^{\prime}$. Kahn and Marković [KM12] prove that in a closed hyperbolic $3$-manifold, the set of nearly geodesic surfaces they construct is ubiquitous. Cooper and Futer [CF19] prove that in a cusped hyperbolic $3$-manifold, there exists a set of $K$-quasi-Fuchsian surfaces which is ubiquitous for some $K>1$, while Kahn and Wright [KW21] prove that for any $K>1$, there exists a ubiquitous set of $K$-quasi-Fuchsian surfaces. Al Assal [AA23] proves the existence of asymptotically Fuchsian surfaces, based on [LM15, KW21]. Theorem 1.6 implies the following.
Corollary 1.9.
Let $M$ be a closed hyperbolic $3$-manifold. A sequence of asymptotically Fuchsian surfaces in distinct commensurability classes is ubiquitous.
The proofs of 1.9 and Theorem 1.6 rely on the rigidity of unipotent flow, i.e., theorems of Ratner [Rat91] and Shah [Sha91a], which is very different from [KM12, CF19, KW21] (but 1.9 and Theorem 1.6 do not give the existence of nearly geodesic surfaces). The analogous statement to 1.9 for cusped hyperbolic $3$-manifolds is a corollary of the main result of [AL25].
Although the property of being filling sorts out a subclass of submanifolds of a hyperbolic $3$-manifold, it also reflects certain topological complexity of the ambient manifold. Recall from Thurston [Thu97, Definition 3.8.1] that a geometric $3$-manifold is a Riemannian $3$-manifold supporting one of the eight model geometries. Nathan Dunfield first suggested the existence of $\pi_{1}$-injective filling surfaces (with infinite fundamental groups) in a non hyperbolic geometric $3$-manifold, i.e., the Hantzsche-Wendt manifold.
Theorem 1.10.
If $M$ is a geometric $3$-manifold and admits a $\pi_{1}$-injective filling homotopy class of surfaces, then $M$ is hyperbolic, spherical, Euclidean, or $\mathbb{H}^{2}\times\mathbb{R}$.
In the proof we explicitly construct some filling surfaces; see Figure 5 for the Euclidean manifolds and Figure 6 for the $\mathbb{H}^{2}\times\mathbb{R}$ ones. We also introduce an elementary inequality connecting the volume of a hyperbolic $3$-manifold with the area of its filling surfaces. By 2.7, a least area surface in a filling $[S]$ cuts $M$ into balls and thus naturally $\textnormal{Area}(S)$ and $\textnormal{Vol}(M)$ satisfy the isoperimetric inequality for $M$, which was pointed out by Wu and Xue [WX22]. The hyperbolic space can be characterized as a space satisfying linear isoperimetric inequalities. Let $\textnormal{rank}(M)$ be the smallest number of generators of $\pi_{1}(M)$.
Theorem 1.11.
@@
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling homotopy class of $\pi_{1}$-injective surfaces. Then any representative $S$ satisfies
(1.1) $\textnormal{Area}(S)>\textnormal{Vol}(M),$
@@
(1.2) $\textnormal{rank}(S)>C\cdot\textnormal{rank}(M).$

### v3

@@ Abstract
Let $M$ be a closed hyperbolic $3$-manifold. We consider equivalence class $[S]$ of connected, orientable, closed surfaces immersed in $M$, modulo homotopy and commensurability. A class $[S]$ is filling if for any $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components that are contractible in $M$. We prove that there exist $\epsilon_{0},\mathfrak{g}_{0}>0$ such that excluding finitely many classes of totally geodesic surfaces of genus $\mathfrak{g}_{0}$, all $(1+\epsilon_{0})$-quasi-Fuchsian $[S]$ are filling. As a corollary, embedded $\pi_{1}$-injective non totally geodesic quasi-Fuchsian surfaces in $M$ have constants bounded below by $1+\epsilon_{0}$. This also gives a curvature gap theorem for embedded minimal surfaces. Each of these filling surfaces separates any pair of distinct points at $\partial\widetilde{M}$. Crucial tools include the rigidity results of Mozes-Shah, Ratner, and Shah. This work is inspired by a question of Wu and Xue whether random geodesics on random hyperbolic surfaces are filling.
@@ 1 Introduction
Let $M$ be a closed hyperbolic $3$-manifold and $\widetilde{M}$ the universal cover. A surface $S$ in $M$ refers to a smooth $\pi_{1}$-injective immersion $S\hookrightarrow M$ of a connected, orientable, closed surface $S$. We denote by $[S]$ an equivalence class of surfaces in $M$ modulo homotopy and commensurability. The genus $\mathfrak{g}[S]$ of $[S]$ is the smallest genus of an orientable surface in $[S]$. A surface $[S]$ is filling if for every $S^{\prime}\in[S]$, $M-S^{\prime}$ consists of components contractible in $M$ (but these components as topological spaces themselves may not be contractible). Rubinstein-Sageev [RS99] define $[S]$ to be strongly filling if for any $p\neq q\in\partial\widetilde{M}$, there is an $\widetilde{S}\subset\widetilde{M}$ covering $S$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S}$. [RS99] proves that a strongly filling class is filling. Our first theorem is the following.
@@
Let $M$ be a closed hyperbolic $3$-manifold. There exist $\mathfrak{g}_{0},\epsilon_{0}>0$ such that excluding finitely many classes of totally geodesic surfaces of genus $\mathfrak{g}_{0}$, all $(1+\epsilon_{0})$-quasi-Fuchsian surfaces are strongly filling.
For any $\epsilon>0$, Jeremy Kahn and Vladimir Marković [KM12b] prove that $M$ has infinitely many $(1+\epsilon)$-quasi-Fuchsian surfaces. Marković suggested to the author that most surfaces constructed in [KM12a] should be filling which help formulate Theorem 1.1. If $[S]$ is strongly filling, it has essential intersection with every geodesic in $M$. If $[S]$ is filling, then by taking a least area representative $S$ (whose existence follows from Schoen-Yau [SY79] and Sacks-Uhlenbeck [SU82]), [RS99] proves that MSM-S consists of topological balls.
For the rest of the paper, we set $\mathfrak{g}_{0}$ and $\epsilon_{0}$ as in Theorem 1.1.
@@
Let $M$ be a closed hyperbolic $3$-manifold. The set of embedded $\pi_{1}$-injective quasi-Fuchsian surfaces in $M$ is decomposed into
@@
    at most finitely many totally geodesic surfaces of genus $\mathfrak{g}_{0}$, and
@@
We construct infinitely many closed hyperbolic $3$-manifolds each of which contains infinitely many embedded $\pi_{1}$-injective quasi-Fuchsian surfaces in 3.20. A combination of theorems due to Thurston [Thu78] and Bonahon [Bon86] proves that a $\pi_{1}$-injective surface $S$ in $M$ is either quasi-Fuchsian or virtually fibered ($S$ lifts to a fiber of a finite cover of $M$ that is a surface bundle over $S^{1}$). The supremum principal curvature of a fibered minimal surface $F$ is >1>1 by [Uhl83, HL21] and the Hausdorff dimension of the limit set of $F$ is $2$. Thus 1.2 implies that any sequence of $\pi_{1}$-injective surfaces in $M$ is eventually self-intersecting or uniformly away from being totally geodesic.
@@
Let $M$ be a closed hyperbolic $3$-manifold. If $S\hookrightarrow M$ is an embedded minimal surface, then $S$ is either a totally geodesic surface of genus $\mathfrak{g}_{0}$ or has supreme principal curvature $\geq\frac{\epsilon_{0}}{\epsilon_{0}+2}$.
Figure 1: The filling nature of nearly geodesic surfaces in a closed hyperbolic $3$-manifold (each dot represents a $(1+\epsilon_{0})$-quasi-Fuchsian surface). The constant $C$ is the universal constant coming from Seppi’s curvature estimate [Sep16].
We summarize Theorem 1.1 in Figure 1. This figure is inspired by [KMS23, Figure 1] and justified in Section 2.1.
Each commensurability classes of Fuchsian surfaces corresponds to a $PSL_{2}(\mathbb{R})$-orbit. A sequence of $(1+\epsilon)$-quasi-Fuchsian minimal surfaces is asymptotically geodesic if $\epsilon\rightarrow 0$, e.g., a sequence of non-commensurable Fuchsian surfaces. The proof of Theorem 1.1 uses Theorem 1.4 and 3.17.
Theorem 1.4.
Let $M$ be a closed hyperbolic $3$-manifold. If $S_{j}]$ is asymptotically geodesic, then the Grassmann $2$-plane bundles $\mathcal{G}_{2}S_{j}$ over minimal surfaces $S_{j}$ converge to $\mathcal{G}_{2}M$ in the Hausdorff metric.
Remark 1.5.
Independently, Al Assal and Lowe [AL25] obtain Theorem 1.4 for finite-volume hyperbolic manifolds with different perspectives and applications. In contrast, Al Assal [AA23] shows that the measure limits of $\mathcal{G}_{2}S_{i}$ achieve every convex combination of $SL_{2}(\mathbb{R})$-invariant ergodic measures.
By a slight abuse of language, by a plane we mean an isometric immersion of $\mathbb{H}^{2}$ in $M$ or $\mathbb{H}^{3}$, or its Grassmann $2$-plane bundle in $\mathcal{G}_{2}M$ or $\mathcal{G}_{2}\mathbb{H}^{3}$. In Theorem 1.4, $\mathcal{G}_{2}M$ is the union of uncountably many planes. If, instead, $S_{i}$ is taken to be the covering of a totally geodesic surface $S$ whose degree tends to infinity as $i\rightarrow\infty$, then the limit of $\mathcal{G}_{2}S_{i}$ is the single plane $\mathcal{G}_{2}S$. Thus Theorem 1.4 reflects the “one versus uncountably many planes duality” suggested by Marković to the author.
3.17 then establishes strong fillingness by the following argument. Suppose there exist nontrivial $\gamma_{i}]\in\pi_{1}(M-S_{i})$. Then a subsequence of $(x_{i},v_{i})$ in the unit tangent bundle $T^{1}\gamma_{i}$ of the geodesic $\gamma_{i}$ converges to $(x,v)$ tangent to a geodesic $\gamma$. By Theorem 1.4, there exist $(p_{j},\Pi_{j})\in\mathcal{G}_{2}S_{j}\rightarrow(x,\Pi)\perp(x,v)$. This implies that $S_{j}$ is nearly orthogonal to $\gamma_{j}$ as well. Then $\partial\tilde{S}_{j}$ separates $\partial\widetilde{\gamma_{j}}$ which means $\pi_{1}(M-S_{j})$ cannot contain $\gamma_{j}$. We need to use [Sep16] which controls the norms of the second fundamental forms of $\tilde{S}_{j}$ by $\epsilon_{j}$ and standard compactness theory for minimal surfaces to prove $\tilde{S}_{j}$ graphically converge to a totally geodesic plane $\widetilde{P}$.
The proof of Theorem 1.4 uses crucially [Rat91, Sha91a] which imply a plane in $M$ is either dense or a closed Fuchsian surface, and uses [MS95] which implies that a sequence of Fuchsian surfaces in distinct commensurability classes equidistribute (become uniformly dense in $\mathcal{G}_{2}M$). Since $S_{i}$ are minimal, their Grassmann bundles are connected closed sets, and their Hausdorff limit is a connected closed subset $\mathcal{G}_{2}S$ of $\mathcal{G}_{2}M$. Any finite union of the Grassmann bundles over Fuchsian surfaces is a disconnected closed set. Thus, $\mathcal{G}_{2}S$, which is closed and $PSL_{2}(\mathbb{R})$-invariant, must contain a dense plane in $\mathcal{G}_{2}M$.
Theorem 1.1 is inspired by Theorem 1.6.
Theorem 1.6.
Let $M$ be a closed hyperbolic $n$-manifold. If $S_{i}$ is a sequence of non commensurable totally geodesic hypersurfaces, then the maximal diameter of components of $M-S_{i}$ tends to 0, the number of components tends to infinity, and all but finitely many $S_{i}]$ are strongly filling.
Since $M$ is aspherical, a component of MSM-S of diameter $2\textnormal{inj}(M)$ is necessarily contractible in $M$. By [BFMS21] (for $n\geq 3$) and [MM22] (for $n=3$), if $M$ contains infinitely totally geodesic hypersurfaces, then $M$ is arithmetic of type I.
An outline of the proof of Theorem 1.6 for the dimension $n=3$ is the following. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Suppose there exist components $C_{i}\subset M-S_{i}$ with $\textnormal{diam}(C_{i})\geq 2R$. Since $\pi^{-1}(C_{i})\subset\mathbb{H}^{3}$ consists of convex components, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length =2R=2R. Up to a subsequence, $\gamma_{i}$ converge to a geodesic segment $\gamma$ of length $2R$. 3.9 proves that the subset of $\mathcal{G}_{2}M$ whose points are tangent to the planes transversally intersecting the central half of $\gamma$ with uniformly bounded angles contains an open set of $\mathcal{G}_{2}M$ of uniformly positive measure (depending only on $R$). Thus eventually $S_{i}$ must intersect the central half of $\gamma$ with uniformly bounded angles. This shows that $S_{i}$ must transversally intersect $\gamma_{i}$ for large $i$, a contradiction. If the maximal diameter of every component of $M-S_{i}$ is less than $2\textnormal{inj}(M)$, then $S_{i}$ is filling.
By 3.10, except for finitely many discrete planes (closed totally geodesic surfaces) $P_{1},\cdots,P_{m}$, every plane $P$ in $M$ is strongly filling. [LM23] points out the rate of equidistribution and hence an upper bound on $m$ depends only on the rate of mixing, $\textnormal{Vol}(M)$, and the injectivity radius. It is surprising that the techniques of [KM12b] and hence the distribution of Kahn-Marković nearly geodesic surfaces also rely on these three geometric invariants.
3.6 proves that a filling Fuchsian surface necessarily has some transverse intersections with all geodesics in $M$. Thus all but finitely many planes cut every geodesic of $M$.
One way to visualize Theorem 1.6 and its transition to Theorem 1.1 is the following. Let $D$ be a convex polyhedron fundamental domain for $M$ in $\mathbb{H}^{3}$. The generators of $\pi_{1}\rightarrow\Gamma\leq PSL_{2}(\mathbb{C})$ are given by finitely many side-pairings of $\partial D$, i.e., $\{g\in\Gamma:g\overline{D}\cap\overline{D}\subset\partial\overline{D}\textnormal{ and}\neq\emptyset\}$ (see 2.6). The preimage in $\mathbb{H}^{3}$ of a filling plane cuts $D$ and $\partial D$ into components so that any geodesic that passes through $D$ must intersect one of the lifts transversally.
[…]
@@
We now describe some motivations from and applications to geometric group theory and virtual properties for strong fillingness. We say $[S]$ in $M$ separates two distinct points $p,q\in\partial\widetilde{M}$ if $S^{\prime}\in[S]$ has a cover $\widetilde{S^{\prime}}\subset\widetilde{M}$ such that $p,q$ belong to different components of $\partial\widetilde{M}-\partial\widetilde{S^{\prime}}$. Kahn-Marković [KM12b] construct a collection of nearly geodesic surfaces which is dense in the Grassmannian $2$-plane bundle $\mathcal{G}_{2}M$. Sageev [Sag97, Sag95] pioneered in a simple but powerful construction of a cocompact action of a word-hyperbolic group on a CAT(0) cube complex $X$ that is dual to a system of walls associated to a finite collection of codimension-$1$ quasiconvex subgroups. Building on Sageev and motivated by obtaining a proper action on $X$, Bergeron-Wise [BW12, Theorem 1.4] formulate a boundary criterion and point out that these nearly geodesic surfaces satisfy the criterion [BW12, Corollary 4.2].
Theorem 1.7 (KM12b, BW12).
@@
Let $M$ be a closed hyperbolic $3$-manifold. There exist a finite collection $\{[S_{1}],\cdots,[S_{m}]\}$ of quasi-Fuchsian surfaces such that every pair of distinct points $p,q\in\partial\widetilde{M}$ is separated by some $\pi_{1}(S_{i})$.
Assuming Theorem 1.7, Agol [Ago13] proves that $\pi_{1}(M)$ is virtually special (in the sense of Haglund-Wise [HW08]), and hence virtually fibered by [Ago08]. For other important consequences in geometric group theory and virtual properties of $3$-manifolds, see Section 3.7, [Ago13, HW08, Sag95, Wis21], and the references therein. There is in general no estimates for the number $m$ in Theorem 1.7 to separate every pair of points at infinity. If $[S]$ is quasi-Fuchsian and strongly filling, then $\{[S]\}$ separates every pair of distinct points at $\partial\widetilde{M}$ and thus satisfies the Bergeron-Wise boundary criterion.
@@
In Section 3.7 we point out that a strongly filling $[S]$ gives a cubulation with a single orbit of hyperplanes (whose stabilizer is essential). Non commensurable strongly filling surfaces give rise to cubulations that are distinct under the isometric actions by $\pi_{1}(M)$ and thus their existence answers a question of Fioravanti-Hagen on [FH21, p. 879].
@@
A stronger notion of separation than fillingness is ubiquity. A collection of immersed surfaces in a hyperbolic 33–manifold $M$ is ubiquitous if, for any pair of disjoint and nonasymptotic planes $\Pi,\Pi^{\prime}\subset\mathbb{H}^{3}$, some surface in the collection has an embedded preimage $\tilde{S}\subset\mathbb{H}^{3}$ that separates $\Pi$ from $\Pi^{\prime}$. Kahn and Marković [KM12b] prove that in a closed hyperbolic $3$-manifold, the set of nearly geodesic surfaces they construct is ubiquitous. Cooper and Futer [CF19] prove that in a cusped hyperbolic $3$-manifold, there exists a set of $K$-quasi-Fuchsian surfaces which is ubiquitous for some $K>1$, while Kahn and Wright [KW21] prove that for any $K>1$, there exists a ubiquitous set of $K$-quasi-Fuchsian surfaces.
Corollary 1.9.
Let $M$ be a closed hyperbolic $3$-manifold (resp. $n$-manifold). A sequence of asymptotically geodesic surfaces (totally geodesic hypersurfaces) is ubiquitous.
The proofs of 1.9 and Theorem 1.4 rely on the rigidity of unipotent flow, i.e., theorems of Ratner [Rat91] and Shah [Sha91a], which is very different from the use of geodesic flows in [KM12b, CF19, KW21] (but 1.9 and Theorem 1.4 do not give the existence of nearly geodesic surfaces). The analogous statement to 1.9 for cusped hyperbolic $3$-manifolds is a corollary of the main result of [AL25].
Although the fillingness sorts out a subclass of submanifolds of a hyperbolic $3$-manifold, it also reflects certain topological complexity of the ambient manifold (this principle fails in general). Recall from Thurston [Thu97, Definition 3.8.1] that a geometric $3$-manifold is a Riemannian $3$-manifold supporting one of the eight model geometries. Nathan M. Dunfield first suggested the existence of $\pi_{1}$-injective filling surfaces (with infinite fundamental groups) in a non hyperbolic geometric $3$-manifold, i.e., the Hantzsche-Wendt manifold.
@@
If $M$ is a geometric $3$-manifold and admits a $\pi_{1}$-injective filling class of surfaces, then $M$ is hyperbolic, spherical, Euclidean, or $\mathbb{H}^{2}\times\mathbb{R}$.
In the proof we explicitly construct some filling surfaces; see Figure 5 for the Euclidean manifolds and Figure 6 for the $\mathbb{H}^{2}\times\mathbb{R}$ ones. We also introduce an elementary inequality connecting the volume of a hyperbolic $3$-manifold with the area of its filling surfaces. By 2.7, a least area surface in a filling $[S]$ cuts $M$ into balls and thus naturally $\textnormal{Area}(S)$ and $\textnormal{Vol}(M)$ satisfy the isoperimetric inequality for $M$. The hyperbolic space can be characterized as one satisfying linear isoperimetric inequalities. Let $\textnormal{rank}(M)$ be the smallest number of generators of $\pi_{1}(M)$.
@@
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling class of $\pi_{1}$-injective surfaces. Then any representative $S$ satisfies
$\textnormal{Area}(S)>\textnormal{Vol}(M),$ (1.1)
@@
$\textnormal{rank}(S)>C\cdot\textnormal{rank}(M).$ (1.2)